The sum of Mrs. Black's age and her daughter's age is 56 years. In 8 years, Mrs. Black will be twice as old as her daughter. How old is her daughter now?
Answer by stanbon(75887)

The sum of Mrs. Black's age and her daughter's age is 56 years. In 8 years, Mrs. Black will be twice as old as her daughter. How old is her daughter now?
---------------------
Equations:
m + d = 56
(m+8) = 2(d+8)
------------------
Substitute for "m" and solve for "d":
(56-d+8) = 2d + 16
----
64 - d = 2d + 16
3d = 48
d = 16 (the daughter is 16 yrs old)
=========================================
Cheers,
Stan H.
=============